# Supervised Algorithmic Fairness in Distribution Shifts: A Survey

### v1

@@ 1 Introduction
To enhance the performance of fairness generalization under distribution shifts, in this survey, we thoroughly examine existing supervised fairness-aware machine learning methods and highlight six commonly used approaches in the literature. Methods falling under feature disentanglement Zhao et al. (2023b) and data augmentation Pham et al. (2023) focus on capturing invariance across various domains. Examining causal graphs and paths Yao and Liu (2023) aids in comprehending how the model’s predictions for different sensitive groups are generated, thereby identifying and addressing potential unfair factors. Adjusting weights of instances or features is a pre-processing method Roh et al. (2023) that enhances model performance by rectifying unfairness in the target domain. The the objective of robust optimization Du and Wu (2021) is to minimize the worst-case loss over various subsets of the training set or other well-defined perturbation sets around the data. Regularization-based methods An et al. (2022) can decrease the correlation between representation and sensitive attributes. It can also enable the transfer of fairness across different domains Schumann et al. (2019).
@@ 2.1 Supervised Algorithmic Fairness
=\>= $\mathbb{P}(f_{{\boldsymbol{\theta}},Z\leftarrow z^{\prime}}(U)=y|X=\mathbf{x},Z=z),\forall y\>\text{and}\>z\neq z^{\prime}$
@@ 2.3 Different Types of Distribution Shifts
Concept Shift Iosifidis et al. (2019) G RW
Iosifidis and Ntoutsi (2020) G OT
Demographic Shift Schumann et al. (2019) G RG
Giguere et al. (2022) G OT
@@ 3 Methods
Causal Inference. Causal models have been widely applied in machine learning to address issues related to model fairness. Structural Causal Models (SCMs) provide a means of explaining machine learning model predictions. Analyzing causal graphs and paths helps understand how the model’s predictions for different groups are formed, thereby identifying and addressing potential unfair factors. A typical fairness causal graph comprises three relational expressions: $Z\to X$, $Z\to Y$, and $Z,X\to Y$. In the context of various distribution shifts, this is equivalent to applying do-operator to corresponding variables, where an intervention on $X$ (denoted $do(X)$) serves the purpose of resetting the distribution of $X$. One can apply the do-operator separately to $X$, $Z$, and $Y$ to generate counterfactual samples, thereby calculating training sample influence on the model’s unfairness Yao and Liu (2023). Similarly, one can analyze the causal dependencies among $X$, $Z$, and $Y$ and their impact on the convergence speed of model training when transitioning to a new distribution Lin et al. (2023b). If only the $do(Z)$ operation is performed, the investigation of counterfactual fairness issues can be conducted. By setting $Z$ to a value different from its original state (e.g., transitioning from male to female), changes in $Y$ may be induced through the propagation of causal paths. It aims for the impact of $do(Z)$ on the predicted values to be minimized Han et al. (2023).
Reweighting. The reweighting approach is a commonly used pre-processing method in statistics and machine learning. It is employed to adjust the weights of samples or features to enhance model performance or correct biases in the dataset. These methods typically involve assigning weights to instances from the source domain to represent the overall distribution of the target domain. One related research employs a sample reweighting technique, where the impact of each instance is adjusted using the classical covariate shift instance weighing function to estimate accuracy metrics Kallus and Zhou (2018). This results in the spurious fairness policy being consistently and uniformly more stringent than necessary for the disadvantaged class. To improve fair training under dependence shift, a decoupling framework was introduced, wherein pre-processing is employed to modify the correlation between sensitive attributes and labels Roh et al. (2023). And then in-processing techniques are subsequently utilized to ensure equal weighting of samples within each $(y,z)$-class.
@@
Table 3: Common datasets for fairness under distribution shifts.

### v3

@@ 1 Introduction
To enhance the performance of fairness generalization under distribution shifts, in this survey, we thoroughly examine existing supervised fairness-aware machine learning methods and highlight six commonly used approaches in the literature. Methods falling under feature disentanglement Zhao et al. (2023b) and data augmentation Pham et al. (2023) focus on capturing invariance across various domains. Examining causal graphs and paths Yao and Liu (2023) aids in comprehending how the model’s predictions for different sensitive groups are generated, thereby identifying and addressing potential unfair factors. Adjusting weights of instances or features is a pre-processing method Roh et al. (2023) that enhances model performance by rectifying unfairness in the target domain. The objective of robust optimization Du and Wu (2021) is to minimize the worst-case loss over various subsets of the training set or other well-defined perturbation sets around the data. Regularization-based methods An et al. (2022) can decrease the correlation between representation and sensitive attributes. It can also enable the transfer of fairness across different domains Schumann et al. (2019).
@@ 2.1 Supervised Algorithmic Fairness
$=$ $\mathbb{P}(f_{{\boldsymbol{\theta}},Z\leftarrow z^{\prime}}(U)=y|X=\mathbf{x},Z=z),\forall y\>\text{and}\>z\neq z^{\prime}$
@@ 2.3 Different Types of Distribution Shifts
Concept Iosifidis et al. (2019) G RW
Shift Iosifidis and Ntoutsi (2020) G OT
Demographic Schumann et al. (2019) G RG
Shift Giguere et al. (2022) G OT
@@ 3 Methods
Causal Inference. Causal models have been widely applied in machine learning to address issues related to model fairness. Structural Causal Models (SCMs) provide a means of explaining machine learning model predictions. Analyzing causal graphs and paths helps understand how the model’s predictions for different groups are formed, thereby identifying and addressing potential unfair factors. A typical fairness causal graph comprises three relational expressions: $Z\to X$, $Z\to Y$, and $Z,X\to Y$. In the context of various distribution shifts, this is equivalent to applying do-operator to corresponding variables, where an intervention on $X$ (denoted $do(X)$) serves the purpose of resetting the distribution of $X$. One can apply the do-operator separately to $X$, $Z$, and $Y$ to generate counterfactual samples, thereby calculating training sample influence on the model’s unfairness Yao and Liu (2023). Similarly, one can analyze the causal dependencies among $X$, $Z$, and $Y$ and their impact on the convergence speed of model training when transitioning to a new distribution Lin et al. (2023). If only the $do(Z)$ operation is performed, the investigation of counterfactual fairness issues can be conducted. By setting $Z$ to a value different from its original state (e.g., transitioning from male to female), changes in $Y$ may be induced through the propagation of causal paths. It aims for the impact of $do(Z)$ on the predicted values to be minimized Han et al. (2023).
Reweighting. The reweighting approach is a commonly used pre-processing method in statistics and machine learning. It is employed to adjust the weights of samples or features to enhance model performance or correct biases in the dataset. These methods typically involve assigning weights to instances from the source domain to represent the overall distribution of the target domain. One related research employs a sample reweighting technique, where the impact of each instance is adjusted using the classical covariate shift instance weighing function to estimate accuracy metrics Kallus and Zhou (2018). To improve fair training under dependence shift, a decoupling framework was introduced, wherein pre-processing is employed to modify the correlation between sensitive attributes and labels Roh et al. (2023). And then in-processing techniques are subsequently utilized to ensure equal weighting of samples within each $(y,z)$-class.
@@
Graph Pokecs 134365 265 Li et al. (2024) Cov. province region working field
Bail-Bs 12024 18 Li et al. (2024) Cov. community race bail
Credit-Cs 16009 13 Li et al. (2024) Cov. community age risk
Table 3: Common datasets for fairness under distribution shifts.
@@
4 Datasets and Evaluation
4.1 Datasets
Tabular datasets. UCI Adult Kohavi and others (1996) includes nearly 48,842 adults and generates a binary income label by determining whether an individual’s income exceeds \$$\$50$k. Gender is commonly chosen as the sensitive attribute. COMPAS Dressel and Farid (2018) consists of 6,167 samples, and its objective is to predict an individual’s recidivism based on criminal history. The dataset assigns race as the sensitive attribute. UCI Adult and COMPAS are also employed for fair anomaly detection, where the binary label indicates whether an instance is anomalous or not. German Asuncion and Newman (2007) comprises a collection of 1,000 samples used for credit scoring prediction of whether an individual is a good or bad credit risk based on various financial features. The dataset includes a binary-sensitive attribute indicating the individual’s gender along with other relevant features. New York Stop-and-Frisk (NYSF) Koh et al. (2021) is a real dataset from the policing activities in New York City. It spans multiple years involving 685,724 records. Due to evident racial bias against African Americans, race is considered a sensitive attribute. Data in different cities serve as distinct domains, and various inspection records can be used as labels, such as stop record, weapon possession, or arrest status. NYSF is also utilized for online settings to address concept shift due to its temporal characteristics Iosifidis and Ntoutsi (2020). In this context, gender is the sensitive attribute.
Image datasets. Colored-MNIST (cMNIST) Arjovsky et al. (2019) purposely establishes a correlation between digits and digit colors in the source data but intentionally anti-correlates them in the target one. Rotated-Colored-MNIST (rcMNIST) Zhao et al. (2023b) is extended from the cMNIST, where each digit is associated with different rotated angles representing domains. For fairness concerns, digit colors are the sensitive attribute correlated to labels. Corlored-Corlored-MNIST (ccMNIST) Arjovsky et al. (2019) is created by colorizing digits and the backgrounds of the MNIST dataset. ccMNIST contains three domains characterized by a different digit color with a different correlation between the class label (same as cMNIST) and sensitive attribute (background colors). Waterbirds Sagawa et al. (2019) is created by merging bird images with backgrounds. The label indicates the bird’s breed, while the sensitive attribute corresponds to the background type. Similar to cMNIST, there exists a spurious correlation between the label and sensitive attribute, generating a dependence shift between the source and target domains. FairFace Kärkkäinen and Joo (2019) contains 108,501 facial images across seven racial categories. A deployment for fairness in distribution shifts is to set the racial groups as domains, gender as the sensitive attribute, and age as class label Zhao et al. (2023a). UTKFace Zhang et al. (2017) comprises 23,708 facial images. Each image is annotated with information regarding the subject’s age, gender, and ethnicity. In this dataset, race is considered a sensitive attribute, while gender is the class label.
Text datasets. BOLD Dhamala et al. (2021) comprises 23,679 English text generation prompts designed to assess societal biases in open-ended language generation, specifically applied to sentiment analysis tasks. The prompts are categorized into five domains based on their topics: profession, gender, race, religion, and political ideology. Gender is considered a binary sensitive attribute within this context.
Graph datasets. Pokecs Dai and Wang (2022) comprises 134,563 graphs extracted from a popular social network in Slovakia, organized according to the users’ respective provinces. Different domains within the dataset are composed of users from two major regions within the corresponding provinces, with the region being treated as a sensitive attribute. Bail-Bs dataset, consisting of nodes representing defendants on bail, is derived from commonly used fair-related graphs obtained from Bail, with race being a sensitive attribute. Credit-Cs is derived from Credit, wherein nodes represent credit card users, following the same partitioning scheme as Bail-Bs. The task involves categorizing customers’ credit risks as high or low based on age as the sensitive attribute.
4.2 Evaluation Metrics for Assessing Fairness
Metrics for Group Fairness. Three fundamental metrics are used for evaluating group fairness. Difference of Demographic Parity ($\Delta_{DP}$) Dwork et al. (2012) and Difference of Equalized Odds ($\Delta_{EO}$) Hardt et al. (2016) are similar to the notions presented in Eq. 1, where $\Delta_{DP}$ and $\Delta_{EO}$ take the absolute probability difference in DP and EO. Difference of Equalized Opportunity ($\Delta_{EOp}$ardt et al. (2016) is similar to $\Delta_{EO}$. The difference lies in that $\Delta_{EOp}$ only requires true positive rates (TPRs) across sensitive subgroups.
$\Delta_{EOp}=$ $|\mathbb{P}(f_{\boldsymbol{\theta}}(X)=1|Z=1,Y=1)$ (5)
$-\mathbb{P}(f_{\boldsymbol{\theta}}(X)=1|Z=0,Y=1)|$
A value of zero for these metrics indicates fair predictions in a target domain.
Metrics for Individual Fairness. Due to the need to measure the similarity between samples, there are relatively few relevant metrics, and consistency Zemel et al. (2013) stands out as one of the more classical ones. Given an input $\mathbf{x}$ to its $k$-nearest neighbors, $kNN(\mathbf{x})$, consistency is formulated:
$\text{consistency}=1-\frac{1}{Nk}\sum_{n}\left|f_{\boldsymbol{\theta}}(\mathbf{x}_{n})-\sum_{j\in kNN(\mathbf{x}_{n})}f_{\boldsymbol{\theta}}(\mathbf{x}_{j})\right|$ (6)
where $N$ represents the total number of samples, and $\mathbf{x}_{n}$ denotes the $n$-th sample. A value close to $1$ represents fairness.
Metrics for Counterfactual Fairness. As counterfactual fairness is a method defined by causal structures, one needs to use the $do$-operator to evaluate it. Assuming $A$ is the intervention target of the $do$-operator, $Y$ is influenced by this intervention. The post-intervention distribution of $Y$ can be expressed as $\mathbb{P}(y_{a})=\mathbb{P}(Y=y|do(A=a))$. Currently, there are primarily two types of metrics.
The Total Causal Effect (TCE) Pearl (2009) of the value change of $Z$ from $z$ to $z^{\prime}$ on $Y=y$ is given by
$\text{TCE}(z,z^{\prime})=|\mathbb{P}(y_{z})-\mathbb{P}(y_{z^{\prime}})|.$ (7)
Given context $O=\mathbf{o}$, the Counterfactual Effect (CE) Shpitser and Pearl (2008) of the value change of $Z$ from $z$ to $z^{\prime}$ on $Y=y$ is given by
$\text{CE}(z,z^{\prime}|\mathbf{o})=|\mathbb{P}(y_{z}|\mathbf{o})-\mathbb{P}(y_{z^{\prime}}|\mathbf{o})|.$ (8)
Smaller TCE and CE indicate that the prediction results are more stable in the counterfactual generation of changing the sensitive attribute, implying greater fairness.
5 Related Research Fields
Unsupervised algorithmic fairness under distribution shifts focuses on addressing unfairness in machine learning models when data labels are not available, and the model needs to adapt to changes in the data distribution. As the currently few unsupervised fairness outlier detection methods, Song et al. Song et al. (2021) leverage deep clustering to discover the intrinsic cluster structure and out-of-structure instances. Meanwhile, adversarial training erases the sensitive pattern for instances of fairness adaptation. Coston et al. Coston et al. (2019) investigate unsupervised domain adaptation for covariate shift between a source and target distribution, highlighting the unavailability of information from the target domain. They employ reweighting to learn weights closely approximating covariate shift weights, defined solely by non-sensitive attributes available in both domains under statistical parity constraints on the source data.
Fairness-aware outlier detection. Unlike conventional classification tasks, outlier detection aims to identify samples where covariate shifts have occurred. Under the influence of sensitive attributes, samples from certain groups are more prone to being identified as outlier samples. Shekhar et al. Shekhar et al. (2021) decompose the fairness in outlier detection into two issues: DP and EOp. They design regularizations to address both problems. Based on the framework of LOF, FairLOF Deepak and Abraham (2021) attempts to correct for such $k$NN neighborhood distance disparities across object groups defined over sensitive attributes. Similarly, based on LOF and adversarial networks, in contrast to FairLOF, AFLOF Li et al. (2022) learns the optimal representation of the original data while concealing the sensitive attribute in the data.
6 Challenges and Future Directions
Fairness under conditional shift. In fact, it is deemed appropriate to align $\mathbb{P}_{X}$ only when $X$ is the causative factor of $Y$. However, it is plausible that $Y$ serves as the cause of $X$, leading to an impact on $\mathbb{P}_{Y|X}$ when there are shifts in $\mathbb{P}_{X}$. The variation of $\mathbb{P}_{Y|X}$ in an unknown domain is referred to as conditional shift Liu et al. (2021). Consequently, certain domain alignment approaches Liu et al. (2021) have proposed aligning the class-conditional distribution $\mathbb{P}_{X|Y}$, under the assumption that $\mathbb{P}_{Y}$ remains unchanged. We believe fairness issues under conditional shift are a promising area for exploration.
Fairness in out-of-distribution detection. While existing work  Han et al. (2023) ensures the absence of anomalous samples during the training process and guarantees fairness in anomaly detection during testing, out-of-distribution detection also necessitates addressing the classification of in-distribution samples. In other words, fairness concerns in out-of-distribution detection involve simultaneously considering multi-class fairness for in-distribution samples and binary-class fairness for out-of-distribution sample detection. Therefore, in the design of the model, it is imperative to balance fairness considerations for both aspects.
7 Conclusion
Ensuring the generalization of model fairness under distribution shifts across domains has emerged as a crucial research focus in recent years. In this survey, we initially compile a list of diverse distribution shifts, offering a brief discussion on the reasons why a predicted model may fail to adapt to each setting. Additionally, we offer a comprehensive review of existing methods addressing fairness generalization in various distribution shifts. Finally, we identify several challenges and suggest potential avenues for future research.
Acknowledgments
This work is supervised by Chen Zhao and Minglai Shao. This work is supported by the National Natural Science Foundation of China program (NSFC #62272338).